eEgg is considering the purchase of a new distributed network computer system to help handle its warehouse inventories. The system costs $65,000 to purchase and install and $34,000 to operate each year. The system is estimated to be useful for 4 years. Management expects the new system to reduce the cost of managing inventories by $67,000 per year. The firm’s cost of capital (discount rate) is 12%.

Required:

1. What is the net present value (NPV) of the proposed investment under each of the following independent situations? (Use the appropriate present value factors from Appendix C, TABLE 1 and Appendix C, TABLE 2.)

1a. The firm is not yet profitable and therefore pays no income taxes.

1b. The firm is in the 20% income tax bracket and uses straight-line (SLN) depreciation with no salvage value. Assume MACRS rules do not apply.

1c. The firm is in the 20% income tax bracket and uses double-declining-balance (DDB) depreciation with no salvage value. Given a four-year life, the DDB depreciation rate is 50% (i.e., 2 × 25%). In year four, record depreciation expense as the net book value (NBV) of the asset at the start of the year.

2. What is the internal rate of return (IRR) of the proposed investment for situations in requirement 1, parts (a) through (c)? Use the IRR builit-in function in Excel to compute the IRR.

eEgg is considering the purchase of a new distributed network computer system to help handle its warehouse inventories. The sComplete this question by entering your answers in the tabs below. Req 1A Reg 1B Reg 1C Req 2 The firm is in the 20% income tComplete this question by entering your answers in the tabs below. Req 1A Req 1B Req 1C Reg 2 The firm is in the 20% income tAnswer is complete but not entirely correct. Complete this question by entering your answers in the tabs below. Req 1A Reg 1BTABLE 1 Present Value of $1 Periods 4% 5% 6% 7% 8% 9% 10% 11% 12% 13% 14% 15% 20% 25% 30% 1 2 3 4 5 6 7 S 9 10 11 12 0.962 0.TABLE 2 Present Value of Annuity of $1 Periods 4% 5% 6% 7% 8% 9% 10% 11% 12% 13% 14% 15% 20% 25% 30% 1 0.962 0.952 0.943 0.93
Answer #1

1a.

Total Savings p.a. => 67000-34000 = 33000.

Year Cashflow Discount factor@ 12% discounted cashflows
0 -65000 1.000 65000
1-4 33000 3.037 100221
NPV 35221

1b.

Depreciation (SLM) = 65000/4 = 16250.

Year Cashflows Depreciation Pre tax Income tax @ 20% Post taxes Discount factor @12% Discounted cashflows
0 -65000 -65000 1.000 65000
1-4 33000 16250 16750 3350 16250-3350+16750 = 29650 3.037 90047
NPV 25047

1c.

Year Cashflows Depreciation Pre tax Income tax @ 20% Post taxes Discount factor @12% Discounted cashflows
0 -65000 -65000 -65000 1.000 -65000
1 33000 32500 500 100 32900 0.893 29380
2 33000 16250 16750 3350 29650 0.797 23631
3 33000 8125 24875 4975 28025 0.712 19954
4 33000 8125 24875 4975 28025 0.636 17824
NPV 25788

2. IRR

a 35.90%
b 29.30%
c 30.40%